OpenStudy (anonymous):

suppose H(x)=(9-6x)^7. find two functions f and g such that (fog)(x)=H(x). neither function can be the identity function. I've been working on this forever and am getting nowhere. suppose H(x)=(9-6x)^7. find two functions f and g such that (fog)(x)=H(x). neither function can be the identity function. I've been working on this forever and am getting nowhere. @Mathematics

OpenStudy (anonymous):

let g(x) = 9-6x and f(x) = x^7

OpenStudy (anonymous):

so f(g(x)) = f(9-6x) = (9-6x)^7

OpenStudy (anonymous):

wow i was way over thinking that one thank you so much!